# FacLens: Transferable Probe for Foreseeing Non-Factuality in Large Language Models

### v3

@@ Appendix I Demo & Case Study
Bad Case User Question: Who won the Turing Award in 2016? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: The Turing Award for 2016 was awarded to Yoshua Bengio, Geoffrey Hinton, and Yann LeCun for their contributions to the development of deep learning. Note: The LLM’s response is incorrect. The correct answer is Tim Berners-Lee.
@@
Bad Case User Question: Where was the 2006 World Expo held? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: The 2006 World Expo was held in Aichi, Japan, from March 25 to September 25, 2005. Note: The LLM’s response is incorrect. The correct answer is Chiang Mai, Thailand.
@@
Bad Case User Question: Who won the Nobel Prize in Physics in 1903? Prediction from FacLens: The LLM knows the factual answer. Main Response from LLM: The Nobel Prize in Physics was not awarded in 1903. The Nobel Prize in Physics was first awarded in 1901, and the first Nobel Prize ceremony was held in 1904. Therefore, there was no Nobel Prize in Physics awarded in 1903. Note: The LLM’s response is incorrect. The correct answer is Henri Becquerel, Pierre Curie, and Marie Curie.

### v4

@@ Appendix E Demo & Case Study
Main Response from LLM: The Turing Award for 2016 was awarded to Yoshua Bengio, Geoffrey Hinton, and Yann LeCun for their contributions to the development of deep learning.
Note: The LLM’s response is incorrect. The correct answer is Tim Berners-Lee.
Bad Case User Question: Where was the 2006 World Expo held?
Prediction from FacLens: The LLM knows the factual answer.
@@
Main Response from LLM: The 2006 World Expo was held in Aichi, Japan, from March 25 to September 25, 2005.
Note: The LLM’s response is incorrect. The correct answer is Chiang Mai, Thailand.
Bad Case User Question: Who won the Nobel Prize in Physics in 1903?
Prediction from FacLens: The LLM knows the factual answer.
@@
Main Response from LLM: The Nobel Prize in Physics was not awarded in 1903. The Nobel Prize in Physics was first awarded in 1901, and the first Nobel Prize ceremony was held in 1904. Therefore, there was no Nobel Prize in Physics awarded in 1903.
Note: The LLM’s response is incorrect. The correct answer is Henri Becquerel, Pierre Curie, and Marie Curie.